OpenStudy (anonymous):

if x=3+2^1/2 then value of x^2+1/x^2

OpenStudy (anonymous):

is your question this: if x = 3 +2^(1/2) then what is the value of (x^2 + 1)/x^2?

OpenStudy (anonymous):

x = 3 +2^(1/2) then what is the value of x^2 + (1/x^2)?

OpenStudy (anonymous):

Solve for the first equation (plug in your calc), and then substitute values for the second equation.

OpenStudy (anonymous):

i solved.... need confirmation

OpenStudy (anonymous):

holiday homework:P

OpenStudy (anonymous):

I have 19.71182229

OpenStudy (anonymous):

i got 15

OpenStudy (campbell_st):

just substitute and evaluate \[\frac{(3 + \sqrt{2})^2 + 1}{(3 + \sqrt{2})^2}\]

OpenStudy (ajprincess):

(x+(1/x))^2=x^2+2+(1/x^2) (x+(1/x))^2-2=x^2+(1/x^2)

OpenStudy (anonymous):

i got 86.239

OpenStudy (anonymous):

wow u guys r really mathmeticians ...every1 is getting different answer

OpenStudy (campbell_st):

I got 1.05132

OpenStudy (anonymous):

it is 15

OpenStudy (anonymous):

15? how

OpenStudy (campbell_st):

but it depends of the equation is it \[x^2 + \frac{1}{x^2}\] or \[\frac{x^2 + 1}{x^2}\]

OpenStudy (anonymous):

Isn't x=3+2^1/2 = 4.414213562?

OpenStudy (anonymous):

it is root 2

OpenStudy (ajprincess):

i get 19.53660216

OpenStudy (anonymous):

dont take it as 1/2

OpenStudy (campbell_st):

and \[x = 3 +\sqrt{2}\]

OpenStudy (anonymous):

Just substitute that value for every x. (3 + sqrt(2)^2) + 1/(3+sqrt(2)^2)

OpenStudy (anonymous):

then answer is 19.537

OpenStudy (anonymous):

@campbell_st I took the equation as x^2+(1/x^2)

OpenStudy (anonymous):

@gaurav555 Can you rewrite both the equations with the BRACKETS where necessary! Everyone is seeing different equations !!!

OpenStudy (anonymous):

x=3+2√2 then value of x²+1/x²

OpenStudy (anonymous):

1. If it is SOLVE, then : x=3+2√2 x= 5.828427 Therefore, substituting x = 5.828427 : x²+1/x² (5.828427² + 1)/(5.828427²) Ans = 1.02943 2. If it is SIMPLIFY, then : = [ (3 + 2√2)² + 1 ] / [ (3 + 2√2)² ] = 1 + [ 1 / (3 + 2√2)² ] = 1 + [ 1 / (9 + 12√2 + 8) ] = 1 + [ 1 / (17 + 12√2) ] Rationalize : = 1 + [ 1 / (17 + 12√2) ] You get : = 1 + (17 - 12√2) = 18 - 12√2 Ans = 18 - 12√2